OpenStudy (camerongirl2019):

simplify: 3(-13)/[5(-3)-7(-2)]

OpenStudy (camerongirl2019):

the little slash is to represent a division sign

OpenStudy (aihberkhan):

Well we want to do this step by step. First lets solve \(3(-13)\). This is \(-39\)

OpenStudy (aihberkhan):

Now lets do \(5(-3)\). This is \(-15\)

OpenStudy (aihberkhan):

Now, \(7(-2)\). This is \(-14\)

OpenStudy (aihberkhan):

So when we put this together, our equation should become \(-39/-15-(-14)\)

OpenStudy (aihberkhan):

Now lets do \(-15-(-14)\). This is \(-1\)

rebeccaxhawaii (rebeccaxhawaii):

\[\frac{ 3(-13) }{ 5(-3)-7(-2) }\] \[\frac{ -39 }{ -15+14 }\]\[\frac{ -39 }{ -1 }\] therefore answer is 39

OpenStudy (aihberkhan):

So our equation becomes \(-39/-1\)

OpenStudy (aihberkhan):

So our answer is \(39\).

OpenStudy (aihberkhan):

Hope that helped! :) @Camerongirl2019

OpenStudy (camerongirl2019):

very much so thank you so much :)

OpenStudy (aihberkhan):

No problem! :) @Camerongirl2019